OpenStudy (zoeeh127):

I need help understanding how to write an equation from point-slope to standard form

jigglypuff314 (jigglypuff314):

point slope form is \(y - y_1 = m(x - x_1)\) and standard form is \(ax + by = c\)

jigglypuff314 (jigglypuff314):

this might seem a little confusing but all you really have to do is distribute the slope (m) then add/subtract #x to both sides to bring it to the side that y is on

jigglypuff314 (jigglypuff314):

for example if you have y - 2 = 3(x - 1) as point-slope form y - 2 = 3x - 3 distribute the 3 -3x -3x minus 3x from both sides ---------------------- -3x + y - 2 = -3 +2 +2 then add 2 to both sides (to get the variables alone) --------------------- -3x + y = -1 which is now in standard form :) do you see how that worked?

OpenStudy (zoeeh127):

can you give me an example with a negative ordered pair

OpenStudy (zoeeh127):

and also thank you i do understand your explaination

jigglypuff314 (jigglypuff314):

you mean a negative ordered pair such as y + 2 = 3(x + 1) ?

OpenStudy (zoeeh127):

yes but preferably one with only one negative coordinate

jigglypuff314 (jigglypuff314):

y - 8 = -4(x + 3) y - 8 = -4x - 12 4x + y - 8 = -12 4x + y= -4